OpenStudy (anonymous):

square root of 36x=x+8

OpenStudy (anonymous):

hello the same process as the other ex. \[\sqrt{36x}=x+8\] square both sides 36x=\[\left( x+8 \right)^{?}\]

OpenStudy (anonymous):

\[6\sqrt{x}=x+8\] \[36x=x^2+16x+64\] \[x^2-20x+64=0\] \[(x-4)(x-16)=0\] \[x=4\] or \[x=16\]

OpenStudy (smurfy14):

x= 4, 16 square each side to get rid of sqrt first

OpenStudy (anonymous):

and of course once again you have to check

OpenStudy (anonymous):

\[\pm \sqrt{36x}=x+8\] \[36x=x^2+16x+64\]

OpenStudy (smurfy14):

^^^ ya satellites got it

OpenStudy (anonymous):

ok now its just a repeat of the above you guys beat me to it :D